# Sublinear bilipschitz equivalence and the quasiisometric classification of solvable Lie groups

### v2

@@ 3.C. Proof of Corollary B and Theorem C
The way Lemma 3.14 is deduced from [Gra23] is the same the way the QI classification was deduced in [KL97]. One may nevertheless long for a direct way of proving this.
@@
Let $m$, $n$, $m^{\prime}$ and $n^{\prime}$ be nonnegative integers. Let $G$, and $G^{\prime}$ be two real semisimple Lie groups with trivial center and no compact factors. Let $\{H_{i}\}_{1\leqslant i\leqslant m}$ and $\{H^{\prime}_{j}\}_{1\leqslant j\leqslant m^{\prime}}$ be families of diagonal Heintze groups satisfying the strong pointed sphere property. Write $G=KAN$, and $G^{\prime}=K^{\prime}A^{\prime}N^{\prime}$. Form the following solvable groups $P$ and $P^{\prime}$ :
$P=\mathbf{R}^{n}\times AN\times\prod_{i=1}^{m}H_{i}\;\text{and}\ P^{\prime}=\mathbf{R}^{n^{\prime}}\times A^{\prime}N^{\prime}\times\prod_{j=1}^{m^{\prime}}H^{\prime}_{j}.$
Let $S$ and $S^{\prime}$ be completely solvable groups. Assume that $S$ and $S^{\prime}$ are quasiisometric, that $\rho_{1}(S)=P$ and that $\rho_{1}(S^{\prime})=P^{\prime}$. Then $m=m^{\prime}$, $n=n^{\prime}$, $G$ and $G^{\prime}$ are isomorphic, and there is a bijection $\sigma$ of $\{1,\ldots,m\}$ such that $H_{i}$ is $O(\log)$-equivalent to $H_{\sigma(i)}$ for all $i$.
@@
The strong pointed sphere property is known to hold for some Heintze groups with nonabelian nilradicals [Pal22, Remark 9]. So the assumptions of Theorem 3.16 are indeed weaker than those of Theorem C. (See also Remark 4.5 for concrete examples.)
@@
Let $L$ and $L^{\prime}$ be semisimple Lie group with no compact factors such that $X=L/K$ and $Y=L^{\prime}/K^{\prime}$. Set $S=AN$, $S^{\prime}=A^{\prime}N^{\prime}$, $n=n^{\prime}=m=m^{\prime}=0$ and apply the theorem to $S$ and $S^{\prime}$. It follows that $L=G$ and $L^{\prime}=G^{\prime}$ are isomorphic, so that $X$ and $Y$ are pluriisometric. ∎
Proof of Theorem C assuming Theorem 3.16.
Let $S$, $S^{\prime}$, $P$ and $P^{\prime}$ be as in the assumptions of the Theorem C. Then $S$, $S^{\prime}$, $P$ and $P^{\prime}$ also satisfy the assumptions of Theorem 3.16, since the groups $H_{i}$ and $H^{\prime}_{j}$ satisfy the strong pointed sphere property as recalled above. By Theorem 3.16, the groups $\mathbf{R}^{n}\times AN$ and $\mathbf{R}^{n^{\prime}}\times A^{\prime}N^{\prime}$ are isomorphic, while, after possibly reindexing the groups $H^{\prime}_{j}$, $H_{i}$ is $O(\log)$-equivalent to $H^{\prime}_{i}$ for all $i$. Now by the main theorem of [Pal20b], $H_{i}$ and $H^{\prime}_{i}$ are isomorphic for all $i$, and so $P$ and $P^{\prime}$ are isomorphic. ∎
We now come to the proof of Theorem 3.16. Applying Theorem A to $P$ and $P^{\prime}$ we will get $O(\log)$-equivalences between the direct factors of $P$ and $P^{\prime}$. The proof consists in establishing which factors can be paired with one another. While it requires some notation, the reality is quite clear: simple factors of $G$ and $G^{\prime}$ must pair to one another, whilst preserving the $\mathbf{R}$-rank one vs. higher $\mathbf{R}$-rank distinction; the strong pointed sphere property further allows to ensure that the remaining factors, namely the Heintze groups of diagonal type with this property, are not paired with the $\mathbf{R}$-rank one factors which do not have it.
@@
Let $n,m,n^{\prime},m^{\prime},S,S^{\prime}$, $P,P^{\prime}$, $G$ and $G^{\prime}$ be as in the assumptions of Theorem 3.16. Decomposing $G$ and $G^{\prime}$ into products of simple factors, we find that there exists $p$, $p^{\prime}$, $q$ and $q^{\prime}$ so that
@@
It follows from Theorem 3.7(1) that the groups $S$ and $P$ on the one hand, $S^{\prime}$ and $P^{\prime}$ on the other hand, are $O(\log)$-bilipschitz equivalent. By assumption, $S$ and $S^{\prime}$ are quasiisometric, especially they are $O(\log)$-bilipschitz equivalent. So $P$ and $P^{\prime}$ are $O(\log)$-bilipschitz equivalent. Note that the coarse type I factors of $P$, resp. $P^{\prime}$ are the groups $H_{i}$, resp. the groups $H^{\prime}_{j}$, while the coarse type II factors are of $P$, resp. $P^{\prime}$ are the groups $A_{k}N_{k}$ for $k\in\{p+1,\ldots,q\}$, resp. $A^{\prime}_{s}N^{\prime}_{s}$ for $s\in\{p^{\prime}+1,\ldots,p^{\prime}+q^{\prime}\}$, which are respectively quasiisometric to the simple Lie groups $G_{k}$ and $G^{\prime}_{s}$ of higher $\mathbf{R}$-rank.
@@
are isomorphic. We now claim that $\sigma_{\mathrm{I}}(\{1,\ldots,p\})=\{1,\ldots,p^{\prime}\}$, especially $p=p^{\prime}$. If it was not the case, then some $H_{i}$, say $H_{i_{0}}$, with $i_{0}>p$, would be $O(u)$-equivalent to a rank one symmetric space. But then the group $\operatorname{SBE}(H_{i_{0}})^{O(\log)}$ would be transitive on its Gromov boundary. This cannot be, as $H_{i_{0}}$ has the strong pointed sphere property. It now follows from (I) and [Pal20a] that the factors of
@@
are pairwise isomorphic, so that $G_{\mathrm{I}}$ and $G^{\prime}_{\mathrm{I}}$ are isomorphic. Finally, by the claim above $\sigma_{\mathrm{I}}(\{p+1,\ldots,\ell\})=\{p+1,\ldots,\ell\}$ so that $H_{i}$ is $O(\log)$-bilipschitz equivalent to $H^{\prime}_{\sigma(i)}$ for all $i\in\{p+1,\ldots,\ell\}$. ∎
@@ 4. Applications to $5$-dimensional solvable Lie groups
An important quasiisometry invariant is the Dehn function. In a subsequent paper [GP] [^2] The full list of Lie groups and computations can currently be found in Section 5 of the arXiv preprint arXiv:2410.05042., using the work of Cornulier and Tessera [CT17], we compute the Dehn functions of all completely solvable groups of exponential growth up to dimension 5.
Based on these computations, we can summarize the contribution of our strategy to these groups.
First, we extract the simply connected solvable Lie groups $G$ such that $\rho_{1}(G)$ decomposes as a product, and then divide them according to their Dehn function. The result is Table 1. The names of the groups are from the classification found in [Mub63] (see also [PSWZ76]); the Lie group are named $G_{d,i}^{\alpha_{1},\ldots,\alpha_{r}}$ where $d$ is the dimension, $i$ is a positive integer, and $\alpha_{1},\ldots,\alpha_{r}$ are parameters (we keep the parameters in the same order as in [Mub63, PSWZ76], but we sometimes use different letters, in accordance with [GP])).
@@
The next step is to determine which of the factors that appear in the decompositions in the $\rho_{1}$ column admit the assumptions of Theorem C or at the very least Theorem 3.16. We do this in the following Lemma.
@@
We conclude that if a group $G$ appearing in Table 1 has quadratic Dehn function then $\rho_{1}(G)$ admits the hypothesis of Theorem C.
Corollary 4.2 (of Theorem C).
@@
Using Theorem 3.16 and part (3) in Lemma 4.1, we can prove that $\mathbf{R}\times G_{4,9}^{\beta}$ is never quasiisometric to $G_{5,30}^{1}$ or $G_{5,37}$ unless $\beta=1$. This was not possible with Theorem C only.

### v3

@@ 3.C. Proof of Corollary B and Theorem D
The way Lemma 3.14 is deduced from [GRA23] is the same the way the QI classification was deduced in [KL97]. One may nevertheless wish for a direct way of proving this.
@@
Let $m$, $n$, $m^{\prime}$ and $n^{\prime}$ be nonnegative integers. Let $G$, and $G^{\prime}$ be two real semisimple Lie groups with trivial center and no compact factors. Let $\{H_{i}\}_{1\leqslant i\leqslant m}$ and $\{H^{\prime}_{j}\}_{1\leqslant j\leqslant m^{\prime}}$ be families of diagonal Heintze groups satisfying the strong pointed sphere property. Let $R$ and $R^{\prime}$ be simply connected nilpotent Lie groups with dimensions $n$ and $n^{\prime}$ respectively as real Lie groups. Write $G=KAN$, and $G^{\prime}=K^{\prime}A^{\prime}N^{\prime}$. Form the following solvable groups $P$ and $P^{\prime}$ :
$P=R\times AN\times\prod_{i=1}^{m}H_{i}\;\text{and}\ P^{\prime}=R^{\prime}\times A^{\prime}N^{\prime}\times\prod_{j=1}^{m^{\prime}}H^{\prime}_{j}.$
Let $S$ and $S^{\prime}$ be completely solvable groups. Assume that $S$ and $S^{\prime}$ are $O(u)$-bilipschitz equivalent for some subadditive function $u$, that $\rho_{1}(S)=P$ and that $\rho_{1}(S^{\prime})=P^{\prime}$. Then $m=m^{\prime}$, $n=n^{\prime}$, $G$ and $G^{\prime}$ are isomorphic, and there is a bijection $\sigma$ of $\{1,\ldots,m\}$ such that $H_{i}$ is $O(u\vee\log)$-equivalent to $H_{\sigma(i)}$ for all $i$.
@@
The strong pointed sphere property is known to hold for some Heintze groups with nonabelian nilradicals [PAL22, Remark 9]. So the assumptions of Theorem 3.16 are indeed weaker than those of Theorem D (see also Remark 4.5 for concrete examples). Moreover, instead of the Euclidean factors of Theorem D, Theorem 3.16 allows more general $R$ and $R^{\prime}$.
@@
Let $L$ and $L^{\prime}$ be semisimple Lie group with no compact factors such that $X=L/K$ and $Y=L^{\prime}/K^{\prime}$. Set $S=AN$, $S^{\prime}=A^{\prime}N^{\prime}$, $n=n^{\prime}=m=m^{\prime}=0$. By assumption, $S$ and $S^{\prime}$ are sublinear bilipschitz equivalent. By Lemma 2.2, $S$ and $S^{\prime}$ are $O(u)$-bilipschitz equivalent for some subadditive function $u$. Applying the theorem to the pair $(S,S^{\prime})$, it follows that $L=G$ and $L^{\prime}=G^{\prime}$ are isomorphic, so that $X$ and $Y$ are pluriisometric. ∎
Proof of Theorem D assuming Theorem 3.16.
Let $S$, $S^{\prime}$, $P$ and $P^{\prime}$ be as in the assumptions of the Theorem D. Then $S$, $S^{\prime}$, $P$ and $P^{\prime}$ also satisfy the assumptions of Theorem 3.16, since the groups $H_{i}$ and $H^{\prime}_{j}$ satisfy the strong pointed sphere property as recalled above. By Theorem 3.16, the groups $\mathbf{R}^{n}\times AN$ and $\mathbf{R}^{n^{\prime}}\times A^{\prime}N^{\prime}$ are isomorphic, while, after possibly reindexing the groups $H^{\prime}_{j}$, $H_{i}$ is $O(u\vee\log)$-equivalent to $H^{\prime}_{i}$ for all $i$. Now by the main theorem of [PAL20b], $H_{i}$ and $H^{\prime}_{i}$ are isomorphic for all $i$, and so $P$ and $P^{\prime}$ are isomorphic. ∎
We now come to the proof of Theorem 3.16. Applying Theorem A to $P$ and $P^{\prime}$ we will get $O(\log)$-equivalences between the direct factors of $P$ and $P^{\prime}$. The proof consists in establishing which factors can be paired with one another. We prove that simple factors of $G$ and $G^{\prime}$ must pair to one another, whilst preserving the $\mathbf{R}$-rank one vs. high $\mathbf{R}$-rank distinction; the strong pointed sphere property further allows to ensure that the remaining factors, namely the Heintze groups of diagonal type with this property, are not paired with the $\mathbf{R}$-rank one factors which do not have it.
@@
Let $n,m,n^{\prime},m^{\prime},S,S^{\prime}$, $P,P^{\prime}$, $R$, $R^{\prime}$, $G$ and $G^{\prime}$ be as in the assumptions of Theorem 3.16. Decomposing $G$ and $G^{\prime}$ into products of simple factors, we find that there exists $p$, $p^{\prime}$, $q$ and $q^{\prime}$ so that
@@
It follows from Theorem 3.7(1) that the groups $S$ and $P$ on the one hand, $S^{\prime}$ and $P^{\prime}$ on the other hand, are $O(\log)$-bilipschitz equivalent. By assumption, $S$ and $S^{\prime}$ are $O(u)$-bilipschitz equivalent, especially they are $O(u\vee\log)$-bilipschitz equivalent. So $P$ and $P^{\prime}$ are $O(u\vee\log)$-bilipschitz equivalent. Note that the coarse type I factors of $P$, resp. $P^{\prime}$ are the groups $H_{i}$, resp. the groups $H^{\prime}_{j}$, while the coarse type II factors are of $P$, resp. $P^{\prime}$ are the groups $A_{k}N_{k}$ for $k\in\{p+1,\ldots,q\}$, resp. $A^{\prime}_{s}N^{\prime}_{s}$ for $s\in\{p^{\prime}+1,\ldots,p^{\prime}+q^{\prime}\}$, which are respectively quasiisometric to the simple Lie groups $G_{k}$ and $G^{\prime}_{s}$ of high $\mathbf{R}$-rank.
@@
are isomorphic. We now claim that $\sigma_{\mathrm{I}}(\{1,\ldots,p\})=\{1,\ldots,p^{\prime}\}$, especially $p=p^{\prime}$. If it was not the case, then some $H_{i}$, say $H_{i_{0}}$, with $i_{0}>p$, would be $O(u)$-equivalent to a rank one symmetric space. But then the group $\operatorname{SBE}(H_{i_{0}})^{O(u\vee\log)}$ would be transitive on its Gromov boundary. This cannot be, as $H_{i_{0}}$ has the strong pointed sphere property. It now follows from (I) and [PAL20a] that the factors of
@@
are pairwise isomorphic, so that $G_{\mathrm{I}}$ and $G^{\prime}_{\mathrm{I}}$ are isomorphic. Finally, by the claim above $\sigma_{\mathrm{I}}(\{p+1,\ldots,\ell\})=\{p+1,\ldots,\ell\}$ so that $H_{i}$ is $O(u\vee\log)$-bilipschitz equivalent to $H^{\prime}_{\sigma(i)}$ for all $i\in\{p+1,\ldots,\ell\}$. ∎
@@ 4. Applications to $5$-dimensional solvable Lie groups
An important quasiisometry invariant is the Dehn function. In a subsequent paper [13], using the work of Cornulier and Tessera [CT17], we compute the Dehn functions of all completely solvable groups of exponential growth up to dimension 5. Based on these computations, we can summarize the contribution of our strategy to these groups.
First, we extract the simply connected solvable Lie groups $G$ such that $\rho_{1}(G)$ decomposes as a product, and then divide them according to their Dehn function. The result is Table 1. The names of the groups are from the classification found in [MUB63] (see also [PSW+76]); the Lie groups are named $G_{d,i}^{\alpha_{1},\ldots,\alpha_{r}}$ where $d$ is the dimension, $i$ is a positive integer, and $\alpha_{1},\ldots,\alpha_{r}$ are parameters (we keep the parameters in the same order as in [MUB63, PSW+76], but we sometimes use different letters, in accordance with [13])).
@@
The next step is to determine which of the factors that appear in the decompositions in the $\rho_{1}$ column admit the assumptions of Theorem D or at the very least Theorem 3.16. We do this in the following Lemma.
@@
We conclude that if a group $G$ appearing in Table 1 has quadratic Dehn function then $\rho_{1}(G)$ admits the hypothesis of Theorem D.
Corollary 4.2 (of Theorem D).
@@
Using Theorem 3.16 and part (3) in Lemma 4.1, we can prove that $\mathbf{R}\times G_{4,9}^{\beta}$ is never quasiisometric to $G_{5,30}^{1}$ or $G_{5,37}$ unless $\beta=1$. This was not possible with Theorem D only.